4a+6(2a-3)-(a+6)=5(a+7)-19
Answer by cassiroberts(11)

4a+6(2a-3)-(a+6)=5(a+7)-19
this is distrubtive property and then solving for A
4a+12a-18--a-6=5a+35-19
16a-18+a-6=5a+16
17a-12=5a+16
subtract 5 from both sides to get
12a-12=16
add 12 to both sides to get
12a=28
divide both sides by 12 to get
28/12=7/6